OpenStudy (anonymous):

Isolate y in the equation 3x + 2y = 8.

OpenStudy (mathstudent55):

First subtract 3x from both sides. Then divide both sides by 2.

OpenStudy (anonymous):

3x + 2y = 8. 2y= -3x+8 \[y=\frac{1}{2}(-3x+8)\] Thus the y is isolated

OpenStudy (texaschic101):

3x + 2y = 8 subtract 3x from both sides 2y = -3x + 8 now divide by 2 to get y by itself y = -3/2x + 4